OpenStudy (anonymous):

x^2=16 answer is +-4? how?

OpenStudy (the_fizicx99):

Hold on let me draw this,

OpenStudy (anonymous):

4*4=16 (-4)(-4)=16, because two negatives multiplied together is a positive.

OpenStudy (anonymous):

Just like (-1)(-1) = 1

OpenStudy (the_fizicx99):

4^2 = 16 16 = 16 it is right, Now as to why it says +-4 is because if its negative, -4^2 = -4 * -4 so two negatives = a positive which is 16 4 * 4 = 16 So regardless if its negative, when you square it it'll be positive :D

OpenStudy (anonymous):

where did the 4 * 4 come from? or the negatives, or anything im so confused

OpenStudy (the_fizicx99):

4^2 is the same as 4 * 4 It's a copy it self multiplied by it self, Same thing like 7^2 7^2 = 7 * 7 or even 5^9 5 ^ 9 = 5*5*5*5*5*5*5*5*5

OpenStudy (anonymous):

WAIT i think i got it

OpenStudy (the_fizicx99):

:P

OpenStudy (anonymous):

4 times 4 = 4^2. You multiply the base number (4 in this case) by itself the number of times in the exponent Also, if you graph x^2=16, you will get a parabola. In the graph, where y=0, x=-4 and 4

OpenStudy (anonymous):

so its just like a quess? like you're like "oooh well 4^2 is 4*4 so that is equal to 16" So like that? hahaha

OpenStudy (the_fizicx99):

Not a guess, 4^2 is 16 4 * 4 = 16

OpenStudy (anonymous):

well yeah but it starts out as x so you guess 4 because 4 times 4 is 16

OpenStudy (anonymous):

4*4 is the same thing as 4+4+4+4

OpenStudy (the_fizicx99):

No you do the inverse, The inverse of a square, is a square root, So, suppose you wanted to find: x^2 = 81 Square rooting both sides gives me x = 9 9^2 = 81

OpenStudy (the_fizicx99):

A square root is ^ 1/2 A square is ^2 A square root is a number that produces a specified quantity when multiplied by itself.

OpenStudy (the_fizicx99):

Do you understand @AmberLynn127